# Failure of uniform laws of large numbers for subdifferentials and beyond

### v1

@@ Abstract
We provide counterexamples showing that uniform laws of large numbers do not hold for subdifferentials under natural assumptions. Our results apply to random Lipschitz functions and random convex functions with a finite number of smooth pieces. Consequently, they resolve the questions posed by Shapiro and Xu [J. Math. Anal. Appl., 325(2), 2007] in the negative and highlight the obstacles nonsmoothness poses to uniform results.
@@ 1 Introduction
Uniform laws of large numbers (LLNs) for gradients are of fundamental importance in stochastic programming [20, Section 7.2.5], and their validity and rate of convergence are now well understood for random smooth functions; see, e.g., [9, 5]. However, the question of whether similar uniform laws hold for their nonsmooth counterparts has remained open for nearly two decades [19, Remark 2]. Positive results are known only for special function classes [17, 19, 21, 24, 15], for enlarged subdifferentials [18, 19], and for weaker notions of convergence [12, 4, 16]; see also the discussion in [20, Section 7.2.6] and [10, p. 471].
To set the stage, let nonempty $X\subset\mathbb{R}^{d}$ be compact, $\Xi\subset\mathbb{R}^{m}$, and $\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be independent and identically distributed (iid) $\Xi$-valued random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. (Throughout, we use boldface for random variables.) Given $f:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ with measurable $f(\cdot,x)$ for all $x\in\mathbb{R}^{d}$, assume the expectation function $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$ is finite valued and $f(\xi,\cdot)$ is $L(\xi)$-Lipschitz on $X$ for any $\xi\in\Xi$, with measurable $L$ and $\mathbb{E}[L(\bm{\xi})]<\infty$. A general uniform LLN for enlarged set-valued mappings appears in [19, Theorem 2] and can be applied to the Clarke subdifferential $\partial_{x}f(\xi,x)=\partial(f(\xi,\cdot))(x)$. Specifically, for any fixed $r>0$, $\mathbb{P}$-almost surely, one has
@@
where $\operatorname{exs}(\cdot;\cdot)$ is the excess of a set relative to another one and the integral is taken in the Aumann sense; see the end of this section for a summary of notation. When $f(\xi,\cdot)$ is further assumed to be subdifferentially regular [14, Definition 7.25] on $X$ for all $\xi\in\Xi$, we have
@@
for any $\xi^{1},\ldots,\xi^{\nu}\in\Xi$ and $x\in X$; see [3, Theorem 2.7.2]. It is then natural to ask, as also explicitly mentioned in [19, Remark 2], whether we can set $r=0$ in (1) and/or (2), which, if both are true, would imply the following uniform law for subdifferentials in the Hausdorff sense:
@@
While the general questions about random Lipschitz functions remain elusive, the uniform law in (3) is actually achievable for several special function classes. One sufficient condition is that, for every $x\in X$, the subdifferential $\partial_{x}f(\xi,x)$ is a singleton for almost every $\xi\in\Xi$, which implies (3) and also that $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$ is continuously differentiable; see [17, Proposition 2.2], [19, Theorem 6], and [20, Theorem 7.52]. Beyond requiring the expectation function to be smooth, the work [21] guarantees (3) when the subdifferential mapping $\partial_{x}f$ satisfies a separable range condition; see [21, Theorem 4] and also [10, Theorem 5.1.31]. While this may be appealing for certain applications, the range of $\partial_{x}f$ is not separable even for the simple random convex function $(\xi,x)\mapsto f(\xi,x)=\max\{x-\xi,0\}$ with $X=[0,1]$ and $\bm{\xi}^{i}$ uniformly distributed on $\Xi=[0,1]$ as pointed out in [12, Example 3.4]; see also [10, Example 5.1.33]. The recent work [15] establishes the uniform law (3) for a subclass of random convex-composite functions. The uniform law in [15, Theorem 5] requires the outer function to be convex, univariate, and deterministic, while the inner function can be random but smooth and Vapnik–Chervonenkis-major (VC-major) [22, Section 2.6.4, Exercise 2.6.13]. It remains open whether the VC-major assumption can be dispensed with. Resolving these questions would deepen our understanding of sample-average approximation (SAA) from stochastic programming and empirical risk minimization in machine learning.
In this paper, we give negative answers to all these questions by providing explicit counterexamples. Our negative results hold under natural assumptions, requiring only simple nonsmooth components. They highlight the obstacles that nonsmoothness poses to uniform results. We report the main results in Sections 2.1 and 2.2, followed by a discussion in Section 2.3. All deferred proofs are collected in Section 3.
@@
We use mostly standard notation. The Borel $\sigma$-algebra on $\Xi$ is denoted by $\mathcal{B}(\Xi)$. A random function $f:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ is Carathéodory if $f(\xi,\cdot)$ is continuous for each $\xi$ and $f(\cdot,x)$ is (Borel) measurable for each $x$; whenever defined, its gradient and Clarke subdifferential at $y$ are denoted by $\nabla_{x}f(\xi,y)=\nabla(f(\xi,\cdot))(y)$ and $\partial_{x}f(\xi,y)=\partial(f(\xi,\cdot))(y)$, respectively. Let $\|\cdot\|$ be the Euclidean norm and let $\mathbb{B}(x,r)=\{y\mid\|y-x\|\leq r\}$. For $\nu\in\mathbb{N}$, we write $\nu]$ for $\{1,\ldots,\nu\}$. A function $g:\mathbb{R}^{d}\to\mathbb{R}$ is $L$-Lipschitz on $X$ (or simply $L$-Lipschitz when $X=\mathbb{R}^{d}$) if $|g(x)-g(y)|\leq L\|x-y\|$ for all $x,y\in X$; it is $L$-smooth on $X$ if it is continuously differentiable ($C^{1}$) with $L$-Lipschitz $\nabla g$ on $X$. A set-valued $S:\Xi\rightrightarrows\mathbb{R}^{d}$ is measurable if $\{\xi\in\Xi\mid S(\xi)\cap C\neq\emptyset\}\in\mathcal{B}(\Xi)$ for every open $C\subset\mathbb{R}^{d}$; its expectation $\mathbb{E}[S(\bm{\xi})]\subset\mathbb{R}^{d}$ is defined in the Aumann sense; see [20, Definition 7.39]. For nonempty sets $C,D\subset\mathbb{R}^{d}$ and a point $x\in\mathbb{R}^{d}$, the distance between $x$ and $C$ is $\operatorname{dist}(x,C)=\inf_{z\in C}\|x-z\|$; the excess of $C$ relative to $D$ is $\operatorname{exs}(C;D)=\mathop{\rm sup}\nolimits_{z\in C}\operatorname{dist}(z,D);$ and the Hausdorff distance between $C$ and $D$ is ${d\kern-1.49994ptl}(C,D)=\max\{\operatorname{exs}(C;D),\operatorname{exs}(D;C)\}$. The $\{0,1\}$-indicator function of a set $C$ is defined by $\mathds{1}_{C}(x)=1$ if $x\in C$ and $\mathds{1}_{C}(x)=0$ otherwise. For $\varepsilon\geq 0$, the $\varepsilon$-subdifferential of a convex function $h:\mathbb{R}^{d}\to\mathbb{R}$ is defined as $\partial^{\varepsilon}h(x)=\{s\in\mathbb{R}^{d}\mid h(y)-h(x)\geq s^{\top}(y-x)-\varepsilon,\forall y\},$ which coincides with the usual convex subdifferential when $\varepsilon=0$.
@@ 2.1 Random Lipschitz Functions
Let $X=[0,1]\subset\mathbb{R}$ and $\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$, each uniformly distributed on $\Xi=[0,1]$. There exist a Carathéodory $f:\Xi\times\mathbb{R}\to\mathbb{R}$ and $\delta^{\nu}\in(0,1]\downarrow 0$ such that the following hold.
@@
The definition of the set $D^{\nu}$ requires that, for every $x\in D^{\nu}$, the function $f(\xi,\cdot)$ is $C^{1}$ on $\mathbb{B}(x,\delta^{\nu})$ for all $\xi\in\Xi$, so that $\nabla_{x}f(\xi,\cdot)$ is well defined (hence measurable) and continuous on $\mathbb{B}(x,\delta^{\nu})$. This is a stringent requirement, and the fact that we can show $D^{\nu}\neq\emptyset$ in (c) and restrict to $D^{\nu}$ in (d) only strengthens the theorem. The following corollary is immediate from Theorem 1.
@@
Under the assumptions of Theorem 1, and for $X$, $\Xi$, $\{\bm{\xi}^{\nu}\}_{\nu}$, $f$, and $\{\delta^{\nu}\}_{\nu}$ constructed there, the following hold.
@@
    For any $\xi\in\Xi$, $\partial_{x}f(\xi,\cdot)$ is outer semicontinuous with $\operatorname{dist}(0,\partial_{x}f(\xi,x))\leq 1$ for all $x\in\mathbb{R}$.
@@
    For any sequence $\{r^{\nu}\in[0,\delta^{\nu}]\}_{\nu}$, $\mathbb{P}$-almost surely, one has
    $\mathop{\rm liminf}_{\nu\to\infty}\sup_{x\in X}\operatorname{exs}\Big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x);\bigcup\nolimits_{y\in\mathbb{B}(x,r^{\nu})}\mathbb{E}[\partial_{x}f(\bm{\xi},y)]\Big)\geq\tfrac{1}{2},$ (4)
    $\mathop{\rm liminf}_{\nu\to\infty}\sup_{x\in X}\operatorname{exs}\Big(\mathbb{E}[\partial_{x}f(\bm{\xi},x)];\bigcup\nolimits_{y\in\mathbb{B}(x,r^{\nu})}\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},y)\Big)\geq\tfrac{1}{2}.$ (5)
@@
The outer semicontinuity, boundedness, and measurability are from standard arguments in the beginning of [19, Section 3]. For measurability, see also [11, Lemma 4]. Part (c) follows from the facts that $\partial_{x}f(\xi,\cdot)=\{\nabla_{x}f(\xi,\cdot)\}$ on $\mathbb{B}(x,\delta^{\nu})$ for any $x\in D^{\nu}\subset X$ and $(\Omega,\mathcal{F})$ is $\mathbb{P}$-complete. ∎
In Corollary 1, we show that neither (1) nor (2) can hold when $r=0$, so (3) is in general impossible for random Lipschitz functions. Consequently, this gives negative answers to the questions posed in [19, Remark 2]. Moreover, it shows that the lower bounds in (4) and (5) remain valid even for $r^{\nu}\downarrow 0$ shrinking sufficiently fast, so that a trivial “smoothing” technique cannot circumvent our negative result.
It is worth noting that the restriction to the subset $D^{\nu}\subset X$ in Theorem 1 yields applications far beyond those for the Clarke subdifferential in Corollary 1. For any $\xi\in\Xi$, the function $f(\xi,\cdot)$ is actually smooth on $D^{\nu}$, and most standard notions of subdifferential coincide there; see, e.g., [14, Chapter 8]. Even for the Clarke subdifferential, we may interchange the sum/expectation and the subdifferential operator, replacing $\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x)$ by $\partial\big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\bm{\xi}^{i},\cdot)\big)(x)$ in (4) and $\mathbb{E}[\partial_{x}f(\bm{\xi},x)]$ by $\partial\mathbb{E}[f(\bm{\xi},\cdot)](x)$ in (5), while the lower bounds there still hold. Consequently, the failure of uniform laws is not an artifact of the particular subdifferential employed, but reflects a more intrinsic obstruction that extends to a broad class of local approximation concepts.

### v2

@@ Abstract
We provide counterexamples showing that uniform laws of large numbers do not hold for subdifferentials under natural assumptions. Our constructions are univariate random Lipschitz functions and bivariate random convex functions with two smooth pieces. Consequently, they resolve the questions posed by Shapiro and Xu [J. Math. Anal. Appl., 325 (2007), 1390–1399] in the negative. They also demonstrate the failure of certain graphical and pointwise laws for subdifferentials, revealing fundamental barriers to the consistency of sample-average approximation and subdifferential approximation.
@@ 1 Introduction
Uniform laws of large numbers (LLNs) for gradients are of fundamental importance in stochastic programming [25, Section 7.2.5], and their validity and rates of convergence are now well understood for random smooth functions; see, e.g., [12, 9]. However, the question of whether similar uniform laws hold for their nonsmooth counterparts has remained open for nearly two decades [24, Remark 2]. Positive results are known only for special function classes [22, 24, 26, 30, 20], for enlarged subdifferentials [23, 24], and for weaker notions of convergence [16, 8, 21]; see also the discussion in [25, Section 7.2.6] and [14, p. 471].
To set the stage, let nonempty $X\subset\mathbb{R}^{d}$ be compact, $\Xi\subset\mathbb{R}^{m}$, and $\bm{\xi},\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be independent and identically distributed (iid) $\Xi$-valued random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Throughout, we use boldface for random variables. Given $f:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ with (Borel) measurable $f(\cdot,x)$ for all $x\in\mathbb{R}^{d}$, assume that the expectation function $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$ is finite-valued and $f(\xi,\cdot)$ is $L(\xi)$-Lipschitz on an open neighborhood of $X$ for any $\xi\in\Xi$, with measurable $L$ and $\mathbb{E}[L(\bm{\xi})]<\infty$. A general uniform LLN for enlarged set-valued mappings appears in [24, Theorem 2] and can be applied to the Clarke subdifferential $\partial_{x}f(\xi,x)=\partial(f(\xi,\cdot))(x)$. Specifically, for any fixed $r>0$, $\mathbb{P}$-almost surely ($\mathbb{P}$-a.s.), one has
@@
where $\operatorname{exs}(\cdot;\cdot)$ is the excess of a set relative to another one and the integral is taken in the Aumann sense; see the end of this section for a summary of notation. When $f(\xi,\cdot)$ is further assumed to be subdifferentially regular [19, Definition 7.25] on $X$ for all $\xi\in\Xi$, we have
@@
for any $\xi^{1},\ldots,\xi^{\nu}\in\Xi$ and $x\in X$; see [7, Theorem 2.7.2].
It is then natural to ask, as also explicitly mentioned in [24, Remark 2], whether we can set $r=0$ in (1) and/or (2), which, if both are true, would imply the following uniform law for subdifferentials in the Hausdorff sense:
@@
While the general questions about random Lipschitz functions remain elusive, the uniform law in (3) is actually achievable for several special function classes. One sufficient condition is that, for every $x\in X$, the subdifferential $\partial_{x}f(\xi,x)$ is a singleton for almost every $\xi\in\Xi$, which implies (3) and also that $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$ is continuously differentiable; see [22, Proposition 2.2], [24, Theorem 6], and [25, Theorem 7.52]. Beyond requiring the expectation function to be smooth, the work [26] guarantees (3) when the subdifferential mapping $\partial_{x}f$ satisfies a separable range condition; see [26, Theorem 4] and also [14, Theorem 5.1.31]. While this may be appealing for certain applications, the range of $\partial_{x}f$ is not separable even for the simple random convex function $f(\bm{\xi},x)=\max\{x-\bm{\xi},0\}$ with $X=[0,1]$ and $\bm{\xi}$ uniformly distributed on $\Xi=[0,1]$ as pointed out in [16, Example 3.4]; see also [14, Example 5.1.33]. The recent work [20] establishes the uniform law (3) for a subclass of random convex-composite functions. The uniform law in [20, Theorem 5] requires the outer function to be convex, univariate, and deterministic, while the inner function can be random but smooth and Vapnik–Chervonenkis-major (VC-major) [28, Section 2.6.4]. It remains unclear whether the VC-type assumption can be removed when the explicit rate in [20, Theorem 5] is not required. Resolving these questions would deepen our understanding of sample-average approximation (SAA) from stochastic programming and empirical risk minimization in machine learning.
In this paper, we give negative answers to all these questions by providing explicit counterexamples. Our negative results hold under natural assumptions, requiring only simple nonsmooth components. We report the main results in Sections 2.1 and 2.2, and then discuss in Section 2.3 the implications of our constructions for the consistency of SAA and for subdifferential approximation. All deferred proofs are collected in Section 3.
@@
We use mostly standard notation. The Borel $\sigma$-algebra on $\Xi$ is denoted by $\mathcal{B}(\Xi)$. A random function $f:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ is Carathéodory if $f(\xi,\cdot)$ is continuous for each $\xi$ and $f(\cdot,x)$ is (Borel) measurable for each $x$; whenever defined, its gradient and Clarke subdifferential at $y$ are denoted by $\nabla_{x}f(\xi,y)=\nabla(f(\xi,\cdot))(y)$ and $\partial_{x}f(\xi,y)=\partial(f(\xi,\cdot))(y)$, respectively. Let $\|\cdot\|$ denote the Euclidean norm, and define $\mathbb{B}(x,r)=\{y\mid\|y-x\|\leq r\}$. For $\nu\in\mathbb{N}=\{1,2,\ldots\}$, write $\nu]=\{1,\ldots,\nu\}$. A function $g:\mathbb{R}^{d}\to\mathbb{R}$ is $L$-Lipschitz on $X$ (or simply $L$-Lipschitz when $X=\mathbb{R}^{d}$) if $|g(x)-g(y)|\leq L\|x-y\|$ for all $x,y\in X$; it is $L$-smooth on $X$ if it is continuously differentiable ($C^{1}$) with $L$-Lipschitz $\nabla g$ on $X$ (or on an open set containing $X$). A set-valued mapping $S:\Xi\rightrightarrows\mathbb{R}^{d}$ is measurable if $\{\xi\in\Xi\mid S(\xi)\cap C\neq\emptyset\}\in\mathcal{B}(\Xi)$ for every open $C\subset\mathbb{R}^{d}$; its expectation $\mathbb{E}[S(\bm{\xi})]\subset\mathbb{R}^{d}$ is defined in the Aumann sense; see, e.g., [25, Definition 7.39]. For nonempty sets $C,D\subset\mathbb{R}^{d}$ and a point $x\in\mathbb{R}^{d}$, the distance between $x$ and $C$ is $\operatorname{dist}(x,C)=\inf_{z\in C}\|x-z\|$; the excess of $C$ relative to $D$ is $\operatorname{exs}(C;D)=\mathop{\rm sup}\nolimits_{z\in C}\operatorname{dist}(z,D);$ and the Hausdorff distance between $C$ and $D$ is ${d\kern-1.49994ptl}(C,D)=\max\{\operatorname{exs}(C;D),\operatorname{exs}(D;C)\}$. The $\{0,1\}$-indicator function of a set $C$ is defined by $\mathds{1}_{C}(x)=1$ if $x\in C$ and $\mathds{1}_{C}(x)=0$ otherwise. For $\varepsilon\geq 0$, the $\varepsilon$-subdifferential of a convex function $h:\mathbb{R}^{d}\to\mathbb{R}$ is defined as $\partial^{\varepsilon}h(x)=\{s\in\mathbb{R}^{d}\mid h(y)-h(x)\geq s^{\top}(y-x)-\varepsilon,\forall y\},$ which coincides with the usual convex subdifferential when $\varepsilon=0$. For $t\in\mathbb{R}$, we write $\lceil t\rceil$ and $\lfloor t\rfloor$ for the ceiling and floor of $t$, respectively.
@@ 2.1 Random Lipschitz Functions
Let $X=[0,1]\subset\mathbb{R}$ and $\bm{\xi},\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$, each uniformly distributed on $\Xi=[0,1]$. There exist a Carathéodory function $f:\Xi\times\mathbb{R}\to\mathbb{R}$ and $\delta^{\nu}\in(0,1]\downarrow 0$ such that the following hold.
@@
The definition of the set $D^{\nu}$ requires that, for every $x\in D^{\nu}$, the function $f(\xi,\cdot)$ is $C^{1}$ on $\mathbb{B}(x,\delta^{\nu})$ for all $\xi\in\Xi$, so that $\nabla_{x}f(\xi,\cdot)$ is well defined (hence $\nabla_{x}f(\cdot,y)$ is measurable) and continuous on $\mathbb{B}(x,\delta^{\nu})$. This is a stringent requirement, and the fact that we can show $D^{\nu}\neq\emptyset$ in (c) and restrict $x$ to $D^{\nu}$ in (d) only strengthens the theorem. The following corollary is immediate from Theorem 1.
@@
Under the assumptions of Theorem 1, and for $X$, $\Xi$, $\{\bm{\xi},\bm{\xi}^{\nu}\}_{\nu}$, $f$, and $\{\delta^{\nu}\}_{\nu}$ constructed there, the following hold.
@@
    For any $\xi\in\Xi$, $\partial_{x}f(\xi,\cdot)$ is outer semicontinuous with $\sup_{z\in\partial_{x}f(\xi,x)}\|z\|\leq 1$ for all $x\in\mathbb{R}$.
@@
    For any sequence $\{r^{\nu}\in[0,\delta^{\nu}]\}_{\nu}$, $\mathbb{P}$-a.s., one has
    $\mathop{\rm liminf}_{\nu\to\infty}\sup_{x\in X}\operatorname{exs}\Big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x);\bigcup\nolimits_{y\in\mathbb{B}(x,r^{\nu})\cap X}\mathbb{E}[\partial_{x}f(\bm{\xi},y)]\Big)\geq\tfrac{1}{2},$ (4)
    $\mathop{\rm liminf}_{\nu\to\infty}\sup_{x\in X}\operatorname{exs}\Big(\mathbb{E}[\partial_{x}f(\bm{\xi},x)];\bigcup\nolimits_{y\in\mathbb{B}(x,r^{\nu})\cap X}\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},y)\Big)\geq\tfrac{1}{2}.$ (5)
@@
The outer semicontinuity, boundedness, and measurability are from standard arguments in the beginning of [24, Section 3]. For measurability, see also [15, Lemma 4]. Part (c) follows from the facts that $\partial_{x}f(\xi,\cdot)=\{\nabla_{x}f(\xi,\cdot)\}$ on $\mathbb{B}(x,\delta^{\nu})$ for any $x\in D^{\nu}\subset X$ and Theorem 1. ∎
In Corollary 1, we show that neither (1) nor (2) can hold when $r=0$, so (3) is in general impossible for random Lipschitz functions. Consequently, this gives negative answers to the questions posed in [24, Remark 2]. Moreover, it shows that the lower bounds in (4) and (5) remain valid even for $r^{\nu}\downarrow 0$ shrinking sufficiently fast, so that a trivial “smoothing” technique cannot circumvent our negative result.
Remark 1.
It is worth noting that the restriction to the subset $D^{\nu}\subset X$ in Theorem 1 yields applications far beyond those for the Clarke subdifferential in Corollary 1. For any $\xi\in\Xi$, the function $f(\xi,\cdot)$ is actually smooth on $D^{\nu}$, and most standard notions of subdifferential coincide there; see, e.g., [19, Chapter 8]. Even for the Clarke subdifferential, we may interchange the sum/expectation and the subdifferential operator, replacing $\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x)$ by $\partial\big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\bm{\xi}^{i},\cdot)\big)(x)$ in (4) and $\mathbb{E}[\partial_{x}f(\bm{\xi},x)]$ by $\partial\mathbb{E}[f(\bm{\xi},\cdot)](x)$ in (5), while the lower bounds there still hold. Consequently, the failure of uniform laws is not an artifact of the particular subdifferential employed, but reflects a more intrinsic obstruction that extends to a broad class of local approximation concepts.